The main attraction at the 1998 international milka chocolate festival in Austria was a giant chocolate bar. The length of the bar was 3.4 feet more than width and the perimeter was 45.8 feet. Find the length and width of the chocolate bar.
Answer by JulietG(1812)

mmmmm...chocolate!


Perimeter = 2 length + 2 width
L = W + 3.4
P = 45.8
Put the known values into the formula
P = 2L + 2W
45.8 = 2(W+3.4) + 2W
Distribute
45.8 = 2W + 6.8 + 2W
Add
45.8 = 4W + 6.8
Subtract 6.8 from each side
39 = 4W
Divide each side by 4
w = 9.75
If the width is 9.75 and the length is 3.4 more than that, the length = 13.15
9.75 + 9.75 + 13.15 + 13.15 = 45.8